OpenStudy (anonymous):

im struggling with these expressions; can anyone help? [(2x^2y)^3/(3xy^3)^2] and [(8x^2)^2/(2y^3)^4]^-1 and [8x^3y^2/4a^2b^4]^-2 please help!!

Directrix (directrix):

@faceman Would you verify that the attached questions are the ones you have entered in your post? Thanks.

OpenStudy (anonymous):

the first one is \[8 x ^{4} \div 9y ^{3}\]

OpenStudy (anonymous):

\[16 y^2 \div 64 x^4\]

OpenStudy (anonymous):

\[16\div (64 \times a^4 \times b^8 \times x^6 \times y^4 )\]